Solve for x: log(x+25)=1+log(2x-7)
Answer by nerdybill(7384)

log(x+25)=1+log(2x-7)
log(x+25) - log(2x-7) = 1
log(x+25)/(2x-7) = 1
(x+25)/(2x-7) = 10^1
(x+25)/(2x-7) = 10
(x+25) = 10(2x-7)
x+25 = 20x-70
25 = 19x-70
95 = 19x
5 = x